OpenStudy (zuha):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y – 4 = 5/2 (x + 3); (–7, 8) Choose the correct answer: y – 8 = - 2/5 (x – 7) y – 8 = 2/5 (x – 7) y – 8 = -2/5 (x + 7) y – 8 = 2/5 (x + 7)

OpenStudy (zuha):

please help

OpenStudy (anonymous):

y – 8 = -2/5 (x + 7) is the Answer because :- As the line is perpendicular, the gradient changes to become -2/5 This perpendicular line passes through the point (-7,8) So -7 and 8 are substituted in the formula :- (y-y1) = m (x-x1) - where m is the gradient of the perpendicular and y1 and x1 are the points through which it passes. This method is also known as one point equation. You can search more about it on Google.

OpenStudy (godlovesme):

@zuha r u there?